### v1

# Beyond the Surface: A Global-Scale Analysis of Visual Stereotypes in Text-to-Image Generation

Abstract
Recent studies have highlighted the issue of stereotypical depictions for people of different identity groups in Text-to-Image (T2I) model generations. However, these existing approaches have several key limitations, including a noticeable lack of coverage of global identity groups in their evaluation, and the range of their associated stereotypes. Additionally, they often lack a critical distinction between inherently visual stereotypes, such as ‘underweight’ or ‘sombrero’, and culturally dependent stereotypes like ‘attractive’ or ‘terrorist’. In this work, we address these limitations with a multifaceted approach that leverages existing textual resources to ground our evaluation of geo-cultural stereotypes in the generated images from T2I models. We employ existing stereotype benchmarks to identify and evaluate visual stereotypes at global scale, spanning 135 nationality based identity groups. We demonstrate that stereotypical attributes are thrice as likely to be present in images of these identities as compared to other attributes. We further investigate how disparately offensive the depictions of generated images are for different nationalities. Finally, through a detailed case study, we reveal how the ‘default’ representations of all identity groups have a stereotypical appearance. Moreover, for the Global South, images across different attributes are visually similar, even when explicitly prompted otherwise. CONTENT WARNING: Some examples may contain offensive stereotypes.
1 Introduction
Figure 1: We identify ’visual’ stereotypes in the generated images of the identity group by grounding the evaluations in existing textual stereotype benchmarks. Yellow boxes denote annotated visual markers of known stereotypes associated with the identity group in the image. We use Stable Diffusion Rombach et al. (2022) to generate images and evaluate them using the stereotypes present in the SeeGULL dataset Jha et al. (2023).
Text-to-Image (T2I) models are increasingly used for generating visual content from textual descriptions, often for downstream tasks such as creative designing, advertising, marketing, and education. More often than not, these T2I models have been trained on large amounts of unverified data from the internet and tend to associate certain words or phrases (either positive or negative) with specific visual concepts. Although, their training enables these models to generate images from simple textual description, it also amplifies and reflects the stereotypes present in the society. Recent studies have highlighted how the generated images for certain identity groups are more stereotypical than others Bianchi et al. (2023a); Qadri et al. (2023); Ghosh and Caliskan (2023). While they notably demonstrate the biases in current state-of-the-art T2I models and how they reinforce existing stereotypes, there are several limitations to their approach.
There is a noticeable lack of coverage of identity groups in the evaluation of stereotypes of the generated images – in terms of the number of global identity groups, their associated stereotypes, and the volume of the images evaluated. This limitation leads to further under-representation of already marginalized minority groups. Moreover, since stereotypes are socially and culturally situated, some stereotypical attributes can be represented more objectively and reliably in an image than others. For example, representation of an ‘attractive’ person, or a ‘terrorist’, is highly subjective and culturally influenced. Existing works on stereotype evaluation often conflate these subjective attributes with more objective visual attributes (like ‘sombrero’, ‘underweight’, etc.,) which can be more objectively identified in an image. Moreover, the current evaluations of the images generated from T2I models blur problematic depictions with stereotypes, and are not necessarily grounded in existing stereotype resources, thereby, impacting the overall stereotype evaluation.
In this work, we address the above limitations by presenting a systematic large-scale cross-cultural evaluation of regional stereotypes present in the generated images from T2I models. We focus on objectively visual stereotypes and evaluate their prevalence in the visual depiction of 135 identity groups across the globe. More specifically, we ground our evaluations in an existing broad-coverage geo-cultural textual dataset, SeeGULL Jha et al. (2023) that incorporates stereotypes for 175 identity groups for the regional dimension. We also study how the default representation of an identity group relate to their stereotypical as well as non-stereotypcal representations. We use Stable Diffusion Rombach et al. (2022) as a representative T2I model for generating images and evaluate over 2000 images across different identity groups. Figure 1 provides a thematic representation of our approach. Although, our methodology uses Stable Diffusion and the SeeGULL repository as a reference, our overall approach is easily extensible to other textual datasets and T2I models. Our main contributions are as follows.
  • We provide a nuanced approach for evaluating Text-to-Image generations for cross-cultural regional stereotypes at scale and with global coverage by leveraging existing resources from textual modality.
  • We make a critical distinction between objectively ‘visual’ and ‘culturally subjective’ stereotypes and identify a list of 385 ‘visual’ attributes that we will release along with their annotations.
  • We introduce a broad-coverage image dataset [^1] For access to the dataset, please fill out the form at : https://forms.gle/3kqsry11D4gLKHQc7 with annotations of visual markers of stereotypes present in the images representing different identity groups.
  • We demonstrate the offensiveness of the generated images and investigate the feasibility of using automated methods employing captioning models to identify visual stereotypes in images at scale.
  • We analyze the representations of several identity groups and demonstrate how T2I models disproportionately lean towards stereotypical representations of identity groups, even when explicitly prompted otherwise.
The above contributions collectively offer a systematic approach for critically examining the stereotypes in images generated from T2I models.
2 Related Work
Stereotypes in Text-to-Image Models
Stable Bias Luccioni et al. (2023) identifies social biases in Text-to-Image systems by analyzing variations in generated images based on gender, ethnicity markers, and professions. They evaluate three popular T2I systems (DALL-E, Stable Diffusion V1.4 and 2), and demonstrate the under-representation of marginalized identities in the generated images. Bianchi et al. (2023b) reveal how T2I models perpetuate dangerous stereotypes related to traits, occupations, and objects using different prompts. They also demonstrate how user efforts to counteract the biases through different prompting strategies prove unsuccessful. Qadri et al. (2023) explore the cultural limitations of T2I models, specifically in South Asian context, and use community-centered perspective to demonstrate how T2I models use an ‘outsiders gaze’ when generating images for South Asian identity groups. Hao et al. (2023) provide a theoretical framework for content moderation for T2I models in their work and propose ways to empirically measure and quantify the harms associated with the generative models. Garcia et al. (2023) partially annotate Google Conceptual Captions dataset with four demographic and two contextual attributes, and demonstrate the demographic biases in image captioning, text-to-image CLIP embeddings, and images generated from T2I models. Naik and Nushi (2023) quantify social biases in occupations, personality traits, and everyday situations, and show that the images generated for default location-neutral prompts resemble those from the United States and Germany. They also demonstrate that not all biases can be mitigated by adding specifications in the input prompt. Basu et al. (2023) quantify the geographical representativeness of the images generated from T2I models. Wang et al. (2023) design a text-to-image association test to measure the implicit stereotypes between concepts and valence, and and their presence in the generated images. Recently, Gustafson et al. (2023) released the FACET dataset containing annotations about person-related attributes and objects such as skin-tine, and measure performance disparity in image classification, object detection and segmentation tasks. Although, there has been a lot of interest in stereotype evaluation recently, we are the first ones to provide a large-scale evaluation study of ‘regional’ stereotypes in T2I models by leveraging existing resources from textual modality. Moreover, we explicitly focus on identifying stereotypes that are primarily visual in nature to reduce the subjectivity in stereotype evaluation.
Stereotype Benchmarks in Textual Modality
There have been several recent efforts to build broad-coverage cross-cultural stereotype resources at scale within textual modalities. For example, StereoSet Nadeem et al. (2021) and CrowS-Pairs Nangia et al. (2020) are stereotype datasets for detecting stereotypes in NLP prediction and classification-based tasks. SeeGULL Jha et al. (2023) is another such stereotype repository which has a broader coverage of global stereotypes containing close to 7000 stereotypes about identity groups spanning 178 countries across 8 different geo-political regions across 6 continents. It also includes stereotypes for state-level identities within the US and India. In this work, we leverage resources from textual modalities, such as SeeGULL, to conduct a comprehensive analysis of stereotypes prevalent in the images generated from Text-to-Image (T2I) models.
3 A Nuanced Analysis of Stereotypes in Images
Stereotypes and their visual depictions are culturally influenced, highly subjective, and vary across regions. For example, in North America, perceptions such as Australians being ‘attractive’, Northern Irish as ‘terrorists’, and Chinese as ‘intelligent’ are quite common Jha et al. (2023). However, these stereotypes are not as common in other countries Jha et al. (2023). The stereotypical attributes lack inherent visual characteristics, and their portrayal is heavily influenced by cultural perceptions. The depiction of a ‘terrorist’, an ‘attractive’ person, or an ‘intelligent’ person, therefore, can vary significantly based on the geographical and cultural context. In contrast, attributes like ‘sombrero’ (often associated with Mexicans), ‘skinny’, ‘underweight’, ‘black’ (stereotypically associated with Sudanese), etc., possess distinct visual markers, allowing a more objective depiction in images (Figure 1). We note here that the terms ‘black’, ‘white’ etc. in this work, as identified in images do not imply race, and instead are about appearance, or appearance-based observed race Hanna et al. (2020), or even the colors themselves.
To ensure a more reliable evaluation of the generated images from Text-to-Image Models for stereotypes, it is crucial to mitigate this subjectivity in visual representation by making a clear distinction between visually identifiable and culturally subjective stereotypes. Therefore, we follow a nuanced two-step approach for evaluating stereotypes in T2I models. As a first step, we answer the fundamental question of which stereotypes can be objectively represented in an image. We identify inherently ‘visual’ attributes which we subsequently use in the second step to detect stereotypes in the generated images by (i) conducting a large-scale annotation study, and (ii) using automated methods. These steps are described in more detail below.
3.1 Identifying Visual Stereotypes
As propounded above, we first undertake a systematic annotation task for annotating ‘visual’ attributes to make a clear distinction between visual and non-visual/subjective attributes. We use an existing stereotype resource in textual modality, SeeGULL Jha et al. (2023), as a reference to ground our evaluations. SeeGULL comprises of (identity, attribute) pairs, where the ‘identity’ term denotes the global identity groups, and ‘attribute’ denotes the associated descriptive attribute terms (adjective/adjective phrase, or a noun/noun phrase), such as (Mexicans, sombrero), (Indians, dark), and (English, stiff upper lip). The dataset has 1994 unique attributes – both visual and non-visual. We conduct an annotation task to separate out the visual attributes from the non-visual ones.
Annotating Visual Nature of Attributes
The annotators are presented with an attribute and a statement of the form of ‘The attribute [attr] can be visually depicted in an image.’ The attribute presented in the sentence is a list of all unique attributes (noun/noun phrases, and adjective/adjective phrases) from SeeGULL. The annotators are required to assign a Likert scale rating to each attribute indicating the extent to which they agree or disagree with the above statement w.r.t. to the given attribute term. The Likert Scale labels were as follows:
  • Strongly Agree: When the attribute can be explicitly identified within an image, like objects, colors, or similar visual elements (e.g., ‘hat,’ ‘sombrero,’ ‘short’).
  • Agree: When the attribute can be deduced from visual cues, albeit not explicitly depicted in the image (e.g., ‘fashionable,’ ‘poor,’ ‘impoverished’).
  • Disagree: When the attribute is challenging to detect visually but may be inferred from visual cues in specific contexts.
  • Strongly Disagree: When the attributes cannot be inferred visually, either explicitly or through visual cues, such as ‘kind,’ ‘talkative,’ ‘warmhearted,’ and the like.
  • Unsure: When annotators are uncertain about the attribute’s visual nature.
Annotator Recruitment
Since the task was in English, we recruited annotators proficient in English reading and writing. The raters were based out of United States. We get annotations for each attribute from 3 annotators that identify with different geographical orgin identities - Asian, European, and North American. Annotators were professional data labelers and were compensated at rates above the prevalent market rates, and respecting the local regulations regarding minimum wage in their respective countries.
Identifying Associated Visual Stereotypes
Using the identified ‘visual’ attributes (please refer to the Appendix 8.1 for results), we filter out the associated ‘visual stereotypes’ in SeeGULL. Figure  2 shows the global distribution of visual attributes and contains examples of visual stereotypes associated with different countries. We observe that Australians (63) have the highest number of visual attributes associated with them followed by Mexicans (46), Indians (34), New Zealand (31), Ethiopians (27), and Japanese (20). The numbers in bracket indicate the number of visual attributes associated with the identity group in SeeGULL. Around 30 identities, including Omani, Ukranian, Swiss, Canadian, Mongolian, etc., just have a single visual attribute associated with them. We use these visual stereotypes for identifying their prevalence in images in the next step.
Figure 2: Global distribution of visual stereotypes across countries. The depth of the color indicates the number of visual stereotypes. A few examples of visual stereotypes of some of the countries have been shown in the figure.
3.2 Detecting Visual Stereotypes in Text-to-Image Generation
More often than not, stereotypes reflected in the images generated using T2I models are implicit associations learned by the models about an identity group during their training. Evaluating these images for the presence of stereotypes is essential so as to not amplify or reinforce the existing bias in the society which can lead to further discrimination of marginalized groups. As a second step to identifying visual stereotypes, we conduct a large-scale annotation task to detect prevalent ‘visual’ stereotypes in the generated images of 135 identity groups across the globe. The scale of our evaluation is significantly larger than the prior work – in terms of image volume, the global coverage of identity groups, and their associated stereotypes.
Generating Images using T2I Models
[…]
Annotating Stereotypes in Images
The annotation task aims to examine whether well-known ‘visual’ stereotypes about various identity groups are reflected in the images generated by Text-to-Image models. We generate 15 images per identity group and present it to the annotators. Each image is accompanied by two sets of attributes: (i) a set of visual attributes stereotypically associated with the identity group (Figure 2), and (ii) an equal number of randomly selected visual non-stereotypical attributes, i.e., attributes that are not stereotypically associated with the identity group. We combine the two sets and randomly show 5 attributes in succession alongside each image, until every attribute has been covered. An additional option of ‘None of the above’ is also displayed with each set. The annotators are asked to select all attribute(s) that they believe are ‘visually’ depicted in the image. Additionally, they are also asked to draw bounding boxes to highlight specific regions, objects, or other indicators that support their selection of the visual attribute within the image. If annotators believe that no attributes are visually represented, they can choose ‘None of the above.’
Annotator Recruitment
Since the task was in English, we recruited annotators proficient in English reading and writing. The raters were based out of United States but identified with different nationalities. We get 3 annotations for each attribute-image pair. Annotators were professional data labelers and were compensated at rates above the prevalent market rates, and respecting the local regulations regarding minimum wage in their respective countries.
Automatically Identifying Stereotypes in Images
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, in addition to annotating stereotypes in images, we also investigate the feasibility of using automated techniques for their detection in images. We use the same set of 15 images per identity group and use an already existing image captioning model, CLIP (Contrastive Language-Image Pre-training) to get top 50 cations or text description of each image.
Building on the already identified visual stereotypes (Section 3), we conduct an analysis of the captions generated for the images of each of the 135 identity groups. We identify the stereotypes in the generated captions by performing a string match. To further understand how uniquely a stereotypical attribute $attr_{s}$ is present in the description of the images of the identity group, we compute a salience score of the attributes w.r.t. the identity group $S(attr_{s},id)$ by using a modified tf-idf metric as follows: $S(attr_{s},id)=tf(attr_{s},id)\cdot idf(attr_{s},C)$ The function $tf(attr_{s},id)$ represents the smoothed relative frequency of attribute $attr_{s}$ for the identity group $id$; and the function $idf(attr_{s},C)$ denotes the inverse document frequency of the attribute term $attr_{s}$ in all the captions ‘C’, reflecting the importance of the attribute across all the captions. A higher salience score of an attribute, indicated a more unique association of the attribute with the identity group. For each identity group, we then extract the most salient visual stereotypes associated with it as present in the captions.
4 Analyzing Stereotypical Depictions in Generated Images
Figure 3: Our approach makes a distinction between "visual" and "non-visual" stereotypes in images. We identify only explicitly present visual stereotypes in the generated images of the identity group.
4.1 Do generated images reflect known stereotypes?
Prior research has shown that generative models have a tendency to exacerbate the bias against minority groups Bianchi et al. (2023b). We uncover commonly observed patterns in the generated images across different identity groups at a large scale by grounding our evaluations in existing textual stereotype resources like SeeGULL. We use the identified visual stereotypes for evaluating and identifying whether already known stereotypical attributes are visually represented in the images of different identity groups. We only consider identity groups that have more than one associated visual stereotypes for our analysis.
We find the likelihood of a stereotypical attribute $attr_{s}$ being present in any image of the identity group $id$, $\mathbb{L}(attr_{s},id)$, as follows. If a stereotypical attribute $attr_{s}$ was shown $n$ times per identity group across all 15 images and selected $k$ times by the annotators, then $\mathbb{L}(attr_{s},id)=\frac{k}{n}$. Figure 3 presents examples of the most likely stereotypes associated with certain identity groups as annotated by different annotators. We observe that the attributes with the highest likelihood for an identity group, also align with the known stereotypes present in SeeGULL. For example, the attributes ‘dark’, ‘thin’, ‘skinny’, and ‘underweight’, the most likely attributes depicting Sudanese individuals, are also known stereotypes in SeeGULL. The attribute ‘poor’ which has a likelihood measure of 0.5 for Bangladeshi, i.e., approximately 50% of images representing Bangladeshis contained a representation of ’poor’, is also an annotated stereotype in SeeGULL. Similarly, attributes ‘beaner’, ‘brown’, ‘festive, ‘dark skin’, and ‘sombrero’ are stereotypes in SeeGULL for Mexicans, and are also present in the images representing a Mexican person. An Indian person is often represented as ‘dark, ‘brown’, and ‘religious’ which are its associated visual stereotypes in SeeGULL.
4.2 Are some identity groups depicted more stereotypically than others?
Although we observe that images of identity groups contain known stereotypical attributes, it is also crucial to understand if representations of some identity groups tend to be more stereotypical than others. To analyze this better, we compute the ‘stereotype tendency’ $\theta_{id}$ for each identity group. We measure the mean likelihood of a stereotype being present in any image representing an identity group $\mathbb{L}({stereo},id)$ and compare it with the likelihood of a randomly selected non-stereotypical attribute being depicted visually in the same set of images $\mathbb{L}({random},id)$.
Let $\mathbb{L}(attr_{s},id)$ denote the likelihood of a stereotypical attribute being present in an image, and $\mathbb{L}(attr_{r},id)$ denote the likelihood of a randomly selected non-stereotypical attribute being present in the image corresponding to an identity group $id$. We select an equal number, $k$, of stereotypical and random attributes for any given identity group for a fair comparison. The likelihood of an image being stereotypical for a given identity group is, then, denoted as $\mathbb{L}({stereo},id)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{L}({attr}_{s}^{i},{id})$, where $i$ denotes the $i$-th stereotypical attribute associated with the identity group. Similarly, the likelihood of a random attribute being present in the image of the identity group is given by $\mathbb{L}({random},id)=\frac{1}{k}\sum_{j=1}^{k}\mathbb{L}({attr}_{r}^{j},{id})$, where $j$ denotes $j$-th attribute selected randomly with the identity group. Computing the ratio between the two for an identity group
$\theta_{id}=\frac{\mathbb{L}(stereo,id)}{\mathbb{L}({random},id)}$
gives us the overall ‘stereotypical tendency’ of the generated images representing an identity group. Higher the ratio for an identity group, greater is the likelihood of a stereotypical representation.
We observe that on average, the visual representation of any identity group is thrice as likely to be stereotypical than non-stereotypical, i.e., the visual stereotypical attributes associated with an identity group are thrice as likely to appear in their visual representation when compared to randomly selected visual non-stereotypical attributes. We also compute $\theta_{id}$ for all 135 identity groups. Table 1 presents the $\theta_{id}$ and the likelihood of images being stereotypical over non-stereotypical for 30 different identity groups. More specifically, images representing Togolese, Zimbabwean, Malian, Guyanese, Iraqi, etc., only contained stereotypical attributes when compared to random attributes. Images representing Nigerians and Sudanese were 27 times more likely to contain stereotypical attributes than randomly selected non-stereotypes. Representations of Ghanaians, Nepalese, and Chinese individuals have an approximately 5 to 6 times higher likelihood of containing stereotypical attributes. We will publicly release the dataset with all the annotated attributes.
Identity $\mathbb{L}$(stereo, id) $\mathbb{L}$(random, id) $\theta_{id}$
Togolese 0.350 0 N/A
Zimbabwean 0.322 0 N/A
Malian 0.289 0 N/A
Guyanese 0.155 0 N/A
Sierra Leonean 0.153 0 N/A
Guatemalan 0.139 0 N/A
Kosovar 0.116 0 N/A
Iraqi 0.114 0 N/A
Swedes 0.100 0 N/A
Danish 0.095 0 N/A
South Sudanese 0.091 0 N/A
Gabonese 0.052 0 N/A
Mauritanian 0.033 0 N/A
Greeks 0.030 0 N/A
Kuwaiti 0.025 0 N/A
Jordanian 0.016 0 N/A
Bhutan 0.016 0 N/A
Moroccan 0.008 0 N/A
Ecuadorian 0.006 0 N/A
Thailand 0.005 0 N/A
Liberian 0.333 0.004 75.00
Panamanian 0.241 0.004 54.37
Lebanese 0.190 0.004 51.53
Mauritian 0.175 0.005 36.83
Sudanese 0.379 0.014 27.35
Nigerian 0.075 0.003 27.29
Libyan 0.243 0.009 25.53
Egyptian 0.127 0.005 24.82
Laotian 0.191 0.009 21.56
Table 1: Likelihood of the representation of an identity group being stereotypical based on the ‘stereotypical tendency’ $\theta_{id}$ for the default representation of the identity group $(id)$.
4.3 Is it feasible to automatically detect stereotypes in images?
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, we also identify the feasibility of using automated methods employing captioning models for stereotype detection. We compare the results with and without using visual stereotypes as a reference. Without visual components to ground the stereotypes, the automated techniques detect a lot of culturally subjective attributes like ‘attractive’, ‘smart’, etc., which are not inherently visual but stereotypically associated with identity groups. However, on using visual attributes as a reference, our approach uncovers only the ‘visual’ stereotypes for an identity group. These stereotypes also have a high likelihood $L(attr_{s},id)$ of being present in the images as marked by the annotators 44.69% of the time. Figure 3 shows the most salient visual stereotypes associated with the identity group. Attributes like ‘sombrero’, ‘dark’, and ‘brown’ were the most salient visual stereotypes for Mexicans; ‘poor’ was highly salient with Bangladeshi, ‘dark’ and ‘thin’ for Sudanese, and ‘elegant’ for French. These attributes were also marked as being present by the annotators. However, our approach also identifies other stereotypical attributes which are not necessarily depicted in the images, e.g., attributes like ‘cow’, and ‘elephant’ for Indians. This could either be a limitation of our approach of detecting stereotypes in the generated captions or could be due to existing errors/biases in the generated captions themselves. Further analysis is required to tease out the errors and biases propagated by different components of such a system.
5 Offensiveness in the Depiction of Identity Groups
Visual representation of some stereotypes can be more offensive than others. For example, the stereotypical depiction of an identity group as ‘poor’ can be considered more offensive when compared to the visual depiction of a stereotype ‘rich’ for another identity group. Therefore, it is also important to understand the offensiveness of these images, and whether certain identity groups have a more offensive representation when compared to others.
The SeeGULL dataset has offensiveness rating for each stereotypical attribute associated with an identity group. We make use of this offensiveness score and infer an overall offensiveness score $O(id)$ for representation of identity groups. Let ${O}(attr_{s},id)$ denote the offensiveness score of a stereotypical attribute $attr_{s}$ associated with an identity group $id$ in SeeGULL. The mean offensiveness score for each identity group is then
$O(id)=\frac{1}{n}\mathbb{L}(stereotype,id)\cdot\sum_{i=1}^{n}{O}(attr_{s}^{i},id)$
where $i$ denotes the $i$-th stereotypical attribute $attr_{s}$ identified as being present in the image representing an identity group by the annotators. Figure 4 visualizes the normalized offensiveness score for different identity groups. We observe that the representations of people from countries in Africa, South America, and South East Asia, are comparatively more offensive. Jordanians, Uruguayans, Gabonese, Laotian, and Albanians have the most offensive representation; whereas Australians, Swedes, Danish, Norwegians, and Nepalese have the least offensive representation.
Figure 4: Offensiveness of the generated images across different countries. The depth of the color increases with the offensive nature of the images.
6 Stereotypical Pull in Generated Images: a Case Study
Figure 5: ‘Stereotypical Pull’: For minority groups, the generative models have a tendency to ‘pull’ the generation of images towards an already known stereotype even when prompted otherwise. The red lines indicate ‘stereotypical’ attributes; the blue lines indicates ‘non-stereotypical attributes’. The numbers indicate the mean cosine similarity score between sets of the image embeddings.
Prior research has shown that Text-to-Image models can have a ‘very homogenized lens’ when representing certain identities Qadri et al. (2023). We conduct a deeper analysis to better understand this across different identity groups and their generated images. If T2I models consistently generate images that align with the stereotypical representations of an identity group, despite being given neutral or non-stereotypical prompts, we term this phenomenon as ‘stereotypical pull.’ This points to the tendency of T2I models to revert to stereotypical depictions, reflecting inherent biases within the models. We illustrate the stereotypical pull across various identity groups using both quantitative and qualitative analysis.
We prompt the T2I models to generate images for 135 identity groups across three different sets by using the following prompts:
  • Default Representation (d): A/An $id$ person; where $id$ denotes the identity group.
  • Stereotypical Representation (s): (i) A/An $id$ person described as $attr_{s}$, (ii) A photo/potrait of a/an $id$ person. described as $attr_{s}$; where $attr_{s}$ is the visual stereotypical attribute associated with the identity group $id$ in SeeGULL.
  • Non-Stereotypical Representation (ns): (i) A/An $id$ person described as $attr_{ns}$, (ii) A photo/potrait of a/an $id$ person described as $attr_{ns}$; where $attr_{ns}$ is a visual attribute not associated with $id$ in SeeGULL.
We generate 15 images per prompt.
Analysing Diversity of Images Globally.
We compute the mean pairwise cosine similarity between the CLIP embedding representations across the above sets of images for identity groups. We use the following three measures.
[…]
Tables  23 and  4 present the mean cosine similarity between the sets of images for all 135 identity groups across all stereotyped and non-stereotyped attributes. We observe that the default representation of any identity group has a higher similarity score with the ‘stereotyped’ images compared to the ‘non-stereotyped’ images for 121 out of 135 identity groups indicating an overall ‘pull’ towards generating stereotypical looking images, irrespective of the identity group. Moreover, for identity groups from global south, the mean similarity score ‘Mean Sim’ across all three measures is constantly higher indicating an overall lack of diversity in their representations even when prompted otherwise.
Disparate Image Diversity for Different Nationalities.
Figure 5 visually demonstrates the ‘stereotypical pull’ for the generated images of the identity groups ‘Bangladeshi’, ‘Swiss’, ‘Ethiopian’, and ‘Australian’ across two dimensions: (i) wealth, and (ii) physical attributes. The default representation of ‘a Bangladeshi person’, looks very similar to the sets of images of ‘a poor Bangladeshi person’ (stereotype), as well as ‘a rich Bangladeshi person’ (non-stereotype). However, the visual representation of ‘a Swiss person’ is quite distinct from the visual representation of ‘a poor Swiss person’ (non-stereotype) and ‘a rich Swiss person’ (stereotype). We observe this pattern even for physical characteristics, where images of ‘an Ethiopian person’, ‘a skinny Ethiopian person (stereotype)’, and ‘a rich Ethiopian person’ (non-stereotype) are all similar looking when compared to that of ‘an Australian person’, ‘a skinny Australian person’ (non-stereotype), and ‘an overweight Australian person’ (stereotype) which are visually quite distinct. The generative models have a tendency to ‘pull’ the image generation towards stereotypes for certain identities even when prompted otherwise.
[…]
Identity Group S(d, s) S(d, ns) S(s, ns) Mean Sim
South Sudanese 0.9824 0.964 0.9609 0.9691
Nigerien 0.9807 0.961 0.9608 0.9675
North Korean 0.9702 0.968 0.956 0.9648
Djiboutian 0.9719 0.9627 0.9596 0.9647
Myanmar 0.9720 0.9603 0.9551 0.9624
Bolivian 0.9688 0.9598 0.9587 0.9624
Mauritanian 0.9736 0.9575 0.9555 0.9622
Malawian 0.9837 0.9514 0.9491 0.9614
Bhutanese 0.9786 0.9546 0.9494 0.9609
Mongolian 0.9719 0.9554 0.9543 0.9605
Table 2: ‘Stereotypical Pull’ observed across generated images of sample identity groups. The default representations of images are more similar to their ‘stereotypical’ representations. Moreover, identity groups from global south consistently have an overall higher mean similarity score S($\cdot$) across their default (d), stereotypical (S), and non-stereotypical (ns) attributes indicating a ‘pull’ towards generating stereotypical images.
7 Conclusion
In this work, we present a comprehensive framework for evaluating Text-to-Image models with a focus on ‘regional stereotypes.’ Leveraging existing stereotype benchmarks in textual resources, we perform a three-fold evaluation of generated images. Firstly, we distinguish between inherently visual stereotypes, which can be represented in an image, and culturally subjective stereotypes. Subsequently, we identify visual stereotypes prevalent in the generated images from T2I models across 135 identity groups by conducting a large-scale annotation task. The annotated images, featuring highlighted visual markers of stereotypes, will be made publicly available for further research. Additionally, we quantify the offensiveness of the generated images across different identity groups, and investigate the feasibility of automatically detecting stereotypes in images. Through an extensive case study, we also demonstrate that generated images for historically marginalized groups exhibit a more visually ‘stereotypical’ appearance, even when Text-to-Image models are explicitly prompted with non-stereotypical attributes. We hope that the presented framework and the dataset will provide valuable insights for understanding and addressing regional stereotypes in visual content generation.
Limitations
Although we cover a wide range of visual stereotypes, annotation about the visual nature of attributes is potentially subjective. While we attempt to account for this subjectivity by (1) using a Likert scale as opposed to binary labels, and (2) capture diversity in our annotator pool w.r.t. gender and geographical region, we recognize that our annotations may still fail to capture other dimensions of subjectivity. Moreover, our evaluation of prevalence of stereotypes in the generated images is limited by the stereotypes present in textual resources like SeeGULL. More regional stereotypes from different textual resources can be included to further expand the overall coverage. We evaluate the images generated using Stable Diffusion V1-4 and believe the results would hold even on other Text-to-Image models, but that needs to be verified in the future work where our analysis may be applied to other generative image models. We chose Stable Diffusion due to its ease of access for academic research, and since our core objective is to demonstrate gaps in existing stereotype bias evaluation approaches, rather than demonstrating biases in any particular or all image generation models. Our work focuses only on regional stereotypes and their presence in the generated images. The proposed framework could be used to experiment with other axes like gender, race, ethnicity, etc., as well. We believe that our approach is extensible to other types of stereotypes, provided such data repositories exist. Future iterations of such data collection and evaluation should take more participatory approach and involve communities with lived experiences on the harms of bias in society.
Ethical Considerations
In this work, we evaluate the stereotypes generated in visual depictions of people, based on their identity associated with a geographical location. We acknowledge that this is a complex notion of identity and will have overlap with other facets of a person’s identity, such as race, which is not a focus of this work. We also emphasize that this work does not cover all possible visual depictions of stereotypes and thus, it’s methods or dataset should not be used as a standard to deem model generations free of stereotypes. This work is an initial step towards a systematic, global-scale evaluations of T2I models for stereotypical depictions, and we hope for future work to build on this to attain more coverage of potential stereotypical and harmful depictions in generations.
Acknowledgements
We thank Utsav Prabhu, and Kathy Meier-Hellstern for their feedback, and advice.